# Giant High-order Nonlinear and Nonreciprocal Electrical Transports Induced by Valley Flipping in Bernal Bilayer Graphene

### v1

@@ IV AC Transport: Valley Oscillation And High-Order Nonlinearity
The critical amplitudes in Fig. 6 have both maximum value $\tilde{E}_{M}=\max(\tilde{E}_{c}^{\alpha})\approx 0.3177E^{*}$ and minimum value $\tilde{E}_{m}=\min(\tilde{E}_{c}^{\alpha})\approx 0.1589E^{*}$ and the angular patterns of the current responses will behave differently for $\tilde{E}<\tilde{E}_{m}$ and $\tilde{E}>\tilde{E}_{M}$. By assuming a $\mathcal{Q}_{1}$-polarized initial state, the second harmonic currents $j^{(2)}$ of the final periodic steady state are calculated and their norms $|j^{(2)}|$ are plotted in Fig. 7 as functions of $\theta$ in polar coordinates. Different rows mean different AC electrical field amplitudes as indicated in the left axis while different columns represent different current components as indicated in the top axis. These plots are colored red and blue according to the sign of their real part $\mathrm{Re}(j^{(2)})$. All the plots in Fig. 7 are antisymmetric about the inversion operation, this arises from the additional tempo-spatial symmetry of the sinusoidal AC electrical field $\tilde{\bm{E}}(t;\theta+\pi)=\tilde{\bm{E}}(t+T/2;\theta)$ as discussed in Sec. IV of the SM. In the first row, the field amplitude is taken as $\tilde{E}=0.05E^{*}<\tilde{E}_{m}$ and no valley flipping occurs. Then the current responses show $C_{3}$-breaking patterns as has already been observed by Lin et al. [27]. In the third row, a much larger electrical field amplitudes $\tilde{E}=0.35E^{*}>\tilde{E}_{M}$ is taken and the broken $C_{3}$ symmetry recovers. This is not surprising since the final states are valley-oscillating states in any electrical field direction as indicated by Fig. 6 and the system will lose its memory about the initial polarization. As discussed in Sec. III of the SM, the critical field strength for valley flipping decreases with hole concentration $n_{h}$, suggesting the $C_{3}$-persevering patterns are more likely to appear near the phase boundary of the 1-pocket state, which was also observed by Lin et al. [27] but might be wrongly interpreted as an unpolarized state in their paper. At the intermediate field amplitudes, for example $\tilde{E}=0.2E^{*}$ in the second row, the angular patterns are similar to the first row in the directions with no valley flipping and similar to the third row in the directions where valley oscillation occurs.

### v2

@@ IV AC Transport: Valley Oscillation And High-Order Nonlinearity
Such non-perturbative behaviors also appear in the I-V curves of the high-harmonic nonlinear responses. In Fig .7(a)(b), the second-harmonic current responses $j^{(2)}$ are plotted as functions of the AC electrical field amplitude $\tilde{E}$. At small electrical field, the currents are well fitted by quadratic functions. However, with the increase of the electrical field amplitude, the I-V curves gradually deviate from the quadratic fits, indicating the failure of the perturbation theory. This agree qualitatively with the experimental results in Ref. [27](FIG. S1). Besides, we also plot the I-V curves of the third-harmonic responses in Fig. 7(c)(d), which show similar behaviors. In spite of the deviation from the perturbation theory, we also find that these I-V curves are discontinuous at the critical electrical field strength. This will help to determine the threshold field experimentally.
Figure 7: (a)(b) I-V curves of the second-harmonic responses. The dashed lines are quadratic fits. (c)(d) I-V curves of the third-harmonic responses. The dashed lines are cubic fits.
@@
The critical amplitudes in Fig. 8 have both maximum value $\tilde{E}_{M}=\max(\tilde{E}_{c}^{\alpha})\approx 0.3177E^{*}$ and minimum value $\tilde{E}_{m}=\min(\tilde{E}_{c}^{\alpha})\approx 0.1589E^{*}$ and the angular patterns of the current responses will behave differently for $\tilde{E}<\tilde{E}_{m}$ and $\tilde{E}>\tilde{E}_{M}$. By assuming a $\mathcal{Q}_{1}$-polarized initial state, the second harmonic currents $j^{(2)}$ of the final periodic steady state are calculated and their norms $|j^{(2)}|$ are plotted in Fig. 9 as functions of $\theta$ in polar coordinates. Different rows mean different AC electrical field amplitudes as indicated in the left axis while different columns represent different current components as indicated in the top axis. These plots are colored red and blue according to the sign of their real part $\mathrm{Re}(j^{(2)})$. All the plots in Fig. 9 are antisymmetric about the inversion operation, this arises from the additional tempo-spatial symmetry of the sinusoidal AC electrical field $\tilde{\bm{E}}(t;\theta+\pi)=\tilde{\bm{E}}(t+T/2;\theta)$ as discussed in Appendix F. In the first row, the field amplitude is taken as $\tilde{E}=0.05E^{*}<\tilde{E}_{m}$ and no valley flipping occurs. Then the current responses show $C_{3}$-breaking patterns. In the third row, a much larger electrical field amplitudes $\tilde{E}=0.35E^{*}>\tilde{E}_{M}$ is taken and the broken $C_{3}$ symmetry recovers. This is not surprising since the final states are valley-oscillating states in any electrical field direction as indicated by Fig. 8 and the system will lose its memory about the initial polarization. At the intermediate field amplitudes, for example $\tilde{E}=0.2E^{*}$ in the second row, the angular patterns are similar to the first row in the directions with no valley flipping and similar to the third row in the directions where valley oscillation occurs.